What are the local maxima, minima, and saddle points of the following function? What are the values of the function at those points? f(x, y) = (1 +xy)(x+y)
1
Expert's answer
2020-10-18T17:16:41-0400

Given, function is


"f(x,y)=(1+xy)(x+y)=x+y+x^2y+y^2x"

Now, critical points are determined by "\nabla f(x,y)=0" , thus


"\nabla f(x,y)=(1+2xy+y^2,1+2xy+x^2)=0\\
\implies 1+2xy+y^2=0,1+2xy+x^2=0\\
\implies y^2=x^2\implies y=\pm x"


Clearly, "y\ne x" as "1+3y^2\ne 0,\forall y\in \mathbb{R}"

Thus, we conclude that "y=-x\implies 1-x^2=0\implies x=\pm1"

Therefore critical points are (1,-1) and (-1,1)

Now, we have to check the Hessian matrix for f at critical points to check whether above critical points are maxima, minima or saddle points.

Now,


"Hf(x,y)=\begin{bmatrix}
 2y&2(x+y)\\
2(x+y)&2x
\end{bmatrix}"

Now, clearly "Hf(1,-1)=\begin{bmatrix}
 -2&0\\0&2
\end{bmatrix}\&Hf(-1,1)=\begin{bmatrix}
 2&0\\0&-2
\end{bmatrix}"

Clearly, both "Hf(1,-1)\&Hf(-1,1)" are diagonal hence their eigen values are "\pm 2"

Hence, both "Hf(1,-1)\&Hf(-1,1)" are semi positive definite ,thus (1,-1) and (-1,1) are saddle points.